# Feature learning as alignment: a structural property of gradient descent in non-linear neural networks

### v3

@@ 4.1 Early time C-NFC dynamics
In order to understand the correlation of centered AGOP and NFM more quantitatively, we will focus on the case of mean-squared error (MSE) loss. Here $\dot{\mathcal{L}}$ corresponds to the diagonal matrix of the residuals $y-f(x)$. If the outputs of the network is 00 on the training data, then ˙$\dot{\mathcal{L}}=Y\equiv{\rm diag}\left(y\right)$, the labels themselves. In that case, the correlation of the time derivatives at initialization is a function of the input-label covariances as well as the PTK matrix:
$\begin{split}&\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)=\,\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\\ &\cdot\text{tr}\left((X^{\top}Y\mathcal{K}YX)^{2}\right)^{-1/2}\cdot\text{tr}\left((X^{\top}Y\mathcal{K}^{2}YX)^{2}\right)^{-1/2}\end{split}$ (6)
In general we expect that these quantities can self-average under the appropriate high-dimensional limits. It is helpful then to write down the average (over initializations) of the covariance. Taking expectation of the first factor,
@@
$\begin{split}&\mathbb{E}_{\theta}\left[\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\right]=\\ &\text{tr}\left(X^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}\right]YXX^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}^{2}\right]YX\right)\\ &+\text{tr}\left(\text{Cov}\left(X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\right)\right)\end{split}$ (7)
A similar decomposition exists for the denominator terms.
The simplest limit is the NTK regime where the width tends to infinity with fixed number of data points and input dimension. Here the PTK matrix will approach its expectation, and the second, covariance term above will tend to 0. This suggests that the first (mean) term encodes much of the interesting phenomenology near initialization for large networks. We use this fact to design a dataset that interpolates between small and large initial derivatives of the C-NFC, even in the non-NTK regime, which will help us demonstrate the validity of our random matrix theory approach.
We focus most of our random matrix theory analysis on a one-hidden layer network, as the calculations quickly become complicated with depth. However, we also show that the mean term can be used to approximately predict the NFC value for Gaussian data with different spectral decay rates for more complicated networks (see Appendix D).
@@ 4.2 Exact predictions with one hidden layer and quadratic activations
We can capture the behavior of both the mean and covariance terms from Equation (7) in certain high dimensional settings. In particular for one hidden layer neural networks with quadratic activations, we can exactly predict the value of Equation (7) in the high dimensional limit. We make the assumption that $X$ and $Y$ are (asymptotically) freely independent of the parameters at initialization, and that the resulting average is close to the value of any specific network initialization (self-averaging). For more details on the assumptions, see Appendix E. We then show numerically that the predicted correlation between the second derivatives of the centered NFM and AGOP matches the observed correlation across individual random instances of the dataset.
To illustrate our analysis, we derive here the numerator of Equation (6). With one hidden-layer, the neural network simplifies to $f(x)=a^{\top}(Wx)^{2}$, where $a$ is the readout layer, and the square is element-wise. Let $M^{(4)}_{X|Y}=(X^{\top}YX)^{2}$, and $M^{(2)}_{X}=X^{\top}X$, and $F_{a}=W^{\top}{\rm diag}\left(a^{2}\right)W$. Then,
@@
We can simplify this alternating product using standard results from random matrix theory, and factor the parameter and data contributions at the cost of increasing the number of terms in the expression (Appendix E).
In the simplified case of isotropic data $X^{\top}X=I$, we have:
$\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\text{tr}\left(M^{(4)}_{X|Y}\right)^{2}\text{tr}\left(F_{a}^{3}\right)+\text{tr}\left(F_{a}\right)\text{tr}\left(F_{a}^{2}\right)\text{tr}\left((\bar{M}^{(4)}_{X|Y})^{2}\right)~{}.$
From these random matrix calculations, we see that the correlations of the derivatives are determined by traces of powers of $M^{(4)}_{X|Y}$ (and $M^{(2)}_{X}$ by the calculations in Section E), which is specific to the dataset, and $F_{a}$, which is specific to the architecture and initialization.
@@
To numerically explore the validity of the random matrix theory calculations, we need a way to generate datasets with different values of $\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)$. We construct a random dataset called the alignment reversing dataset, parameterized by a balance parameter $\gamma\in(0,1]$, to adversarially disrupt the NFA near initialization in the regime that width $k$, input dimension $d$, and dataset size $n$ are all equal ($n=k=d=1024$). By Proposition 8, for the aforementioned neural architecture, the expected second derivative of the centered NFM satisfies, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}\right]YX=(X^{\top}YX)^{2}$, while the expected second derivative of the centered AGOP, $\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}^{2}\right]YX$, has an additional component $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$. Our construction exploits this difference in that $X^{\top}X$ becomes adversarially unaligned to $X^{\top}YX$ as the balance parameter decreases.
@@
Dataset Construction
The data consists of a mixture of two distributions from which two subsets of the data $X^{(1)}$ and $X^{(2)}$ are sampled from, and is parametrized by the balance $\gamma\in(0,1]$ and two variance parameters $\epsilon_{1},\epsilon_{2}>0$. The subset $X^{(1)}$ which has label $y_{1}=1$ and constitutes a $\gamma$ fraction of the entire dataset, is sampled from a multivariate Gaussian distribution with mean 00 and covariance
$\Sigma=\bm{1}\bm{1}^{\top}+\epsilon_{1}\cdot I~{}.$
Then the second subset, $X^{(2)}$, is constructed such that $(X^{(2)})^{\top}X^{(2)}\approx((X^{(1)})^{\top}X^{(1)})^{-2}$, and has labels $y_{2}=0$. Then, for balance parameter $\gamma$ sufficiently small, the AGOP second derivative approximately satisfies,
$\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]$ $\sim X^{\top}YXX^{\top}XX^{\top}YX$
$=(X^{(1)})^{\top}X^{(1)}X^{\top}X(X^{(1)})^{\top}X^{(1)}$
$\approx(X^{(1)})^{\top}X^{(1)}(X^{(2)})^{\top}X^{(2)}(X^{(1)})^{\top}X^{(1)}$
$\approx I~{},$
In our experiment, we sample multiple random datasets with this construction and compute the predicted and observed correlation of the second derivatives of the centered NFC at initialization. For specific details on the construction, see Appendix F.
@@ 5.1 Feature learning and initialization
One factor that modulates the level of feature learning is the scale of the initialization in each layer. In our experiments training networks with unmodified gradient descent, we observe that the centered NFC will increasingly dominate the uncentered NFC with decreasing initialization (third panel, first row, Figure 4). Further, as the centered NFC increases in contribution to the uncentered quantity, the strength of the UC-NFC, and to a lesser extent, the C-NFC increases (Figure 4). The decreases in correspondence between these quantities is also associated with a decrease in the feature quality of the NFM (Appendix H), for the chain monomial task.
For fully-connected networks with homogeneous activation functions, such as ReLU, and no normalization layers, decreasing initialization scale is equivalent to decreasing the scale of the outputs, since we can write $f(Wx)=a^{-p}f(aWx)$ for any scalar $a$ for any homogenous activation $f$. This in turn is equivalent to increasing the scale of the labels. Therefore, decreasing initialization forces the weights to change more in order to fit the labels, leading to more change in $F$ from its initialization. Conceptually, this aligns with the substantial line of empirical and theoretical evidence that increasing initialization scale or output scaling transitions training between the lazy and feature learning regimes [6, 32, 33, 34].
@@ 5.2 Speed Limited Optimization
$W^{(\ell)}_{t+1}\leftarrow W^{(\ell)}_{t}+\eta\cdot C_{\ell}\cdot\frac{\nabla_{W^{(\ell)}_{t}}\mathcal{L}}{\|\nabla_{W^{(\ell)}_{t}}\mathcal{L}\|}$
where the hyperparameter $C_{\ell}\geq 0$ controls the amount of learning in layer $\ell$. We expect this rule to increase the strength of the UC-NFC in layers where $C_{\ell}$ is large relative to $C_{m}$ for $m\neq\ell$, as $W^{(\ell)}$ will be forced to change significantly from initialization. As a result, $\left\|W_{0}^{(\ell)}\right\|\left\|W_{t}^{(\ell)}\right\|^{-1}\rightarrow 0$ for large $t$, causing the centered and uncentered NFC to coincide for this layer.
We demonstrate the effects of SLO on the chain-monomial task (Equation (2)). We found that fixing the learning speed to be high in the first layer and low in the remaining layers causes the ratio of the unnormalized C-NFC to the UC-NFC to become close to $1$ across initialization scales (Figure 4). The same result holds for the SVHN dataset (see Appendix L). We note that this intervention can be applied to target underperforming layers and improve generalization in deeper networks (see Appendix J for details).
We observe that both the C-NFC and the UC-NFC become close to $1$ after training with SLO, independent of initialization scale (Appendix H). Further, the quality of the features learned, measured by the similarity of $F$ and $\bar{G}$ to the true EGOP, significantly improve and are more similar to each other with SLO, even with large initialization. In contrast, in standard training the UC-NFA fails to develop with large initialization, as $F$ resembles identity (no feature learning), while $\bar{G}$ only slightly captures the relevant features.
@@ 6 Discussion
Our theoretical analysis in this work shows that the PTK feature covariance at initialization has a relatively simple structure in terms of the weights and the data. To predict the NFC later in training, we will likely need to account for the change in this matrix. One should be able to predict this development at short times by taking advantage of the fact that eigenvectors of the Hessian change slowly during training [35]. An alternative approach would be to use a quadratic model for neural network dynamics [36, 37] which can capture dynamics of the empirical NTK (and therefore the PTK).
Depth and general activations
We provide an analytic prediction for the C-NFC at initialization for one hidden layer networks with quadratic activation functions, and provide first-order predictions for two hidden layer networks with quadratic and ReLU activations. A followup direction would be to use random matrix analysis to more accurately predict the NFC for deep architectures with general activations. A promising approach in this direction is to approximate non-linear activations with a linear expansion and independent correction terms [38, 39, 40, 41]. Alternatively, we may hope to establish a general spectral condition on $\mathcal{K}$ that facilitates high correlation between $\dot{W}^{\top}\dot{W}$ and $\dot{W}^{\top}K\dot{W}$ at early training times.
@@
There are prior works that implement differential learning rates across layers [42, 43], though these differ from differential learning speeds as in SLO, in which the L2-norm of the gradients are fixed, and not the scaling of the gradients. Our intervention is more similar to the LARS [44] and LAMB [45] optimizers that fix learning speeds to decrease training time in ResNet and BERT architectures. However, these optimizers fix learning speeds to the norm of the weight matrices, while our intervention sets learning speeds to free hyperparameters in order to increase feature learning, as measured by the NFC. We also note as an additional explanation for the success of SLO in our experiments, the PTK feature covariance may change slowly (or just in a single direction) with this optimizer, because the learning rate is small for later layers, allowing the first layer weights to align with the PTK.
@@
We demonstrate in this work that the dataset (by our construction in Section 4) and optimizer (with SLO in Section 5) play a significant role in the strength of feature learning. Important future work would be to understand other settings where significant empirical differences exist between optimization choices. In particular, analyzing the NFC may clarify the role of gradient batch size and adaptive gradient methods in generalization [46].

### v4

@@ 4.2 Maximum C-NFC for data on the sphere
We will consider $f$ trained using mean-squared error (MSE) loss. With this loss, $\dot{\mathcal{L}}$ corresponds to the diagonal matrix of the residuals $y-f(x)$. If the outputs of the network is 00 on the training data, then ˙$\dot{\mathcal{L}}=Y\equiv{\rm diag}\left(y\right)$, the labels themselves. We can guarantee this by either of the following methods:
Method 7 (Subtract copy).
Subtracting off an identical (untrained) copy of the neural network from each output at initialization. I.e. we train the parameters of the network $f$ on the loss with respect to outputs $\hat{f}_{t}(x)=f_{t}(x)-f_{0}(x)$, where $f_{0}(x)$ is a copy of $f$ at initialization and $f_{t}(x)=f(x;a_{t},W_{t})$
Method 8 (Small initialization).
Initializing $W=0$ or $\|a\|=\epsilon$ for $\epsilon>0$ arbitrarily small.
We now state our main theorem.
Theorem 9 (Maximum C-NFC).
Suppose the data $X$ are sampled uniformly at random from $\mathbb{S}^{d-1}$ and the labels are generated by $f^{*}$ satisfying Assumptions 4 and 5. Suppose we train $f$ with MSE loss and initialize the network so that the initial outputs are 00 by either Method 7 or Method 8 above. Assume the activation function $\phi$ satisfies Assumption 6. We consider the regime that $\omega_{d}\left(d\log^{2\ell+1}d\right)\leq n\leq o_{d}\left(d^{2-\delta}\right)$ for some $\delta>0$, and width $k\rightarrow\infty$. Then, at $t=0$ in training (in the setting of Proposition 3), almost surely over the random $X$ as $n,d\rightarrow\infty$,
$\rho\left(\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}\bar{W}),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}(\bar{W}^{\top}K\bar{W})\right)\rightarrow 1~{}.$
The proof follows from the mean term of $\mathcal{K}$ giving the leading order terms in the C-NFC derivatives calculation here, and $\underline{1}\underline{1}^{\top}$ is a rank-$1$ projector. The proof is deferred to Appendix C.
We clarify that although we take infinite width, we are not necessarily in the NTK regime, as we allow for arbitrary scaling of the weights, including the $\mu$P parametrization (Yang and Hu, 2020). Hence, our setting allows for feature learning.
In the next section we will construct a dataset which interpolates between adversarial and aligned eigenstructure to demonstrate the range of possible values the derivative correlations can take, and theoretically predict their values.
4.3 Early time C-NFC dynamics in co-scaling regime
Another interesting limit is the linear co-scaling regime, where $n,d,k\to\infty$ with $n/d\equiv\psi_{1}$ and $k/d\equiv\psi_{2}$. Here, we show it is possible to theoretically predict the correlation of the derivatives of the centered NFM and AGOP.
In this regime, we expect the following two properties (Adlam and Pennington, 2020), which we state as assumptions. First, our key quantities can be written as traces of products of large random matrices; as the dimensions of all such matrices increase, we expect these traces to converge to their average values. This property is known as self-averaging.
Assumption 10 (Self-averaging).
We assume that the expected traces appearing in the NFC across initializations are equal to the traces themselves.
The elements of weight matrices $W$ are drawn from independent Gaussians (i.i.d. within each matrix). If the data $X$ were also standard Gaussian, we would be able to apply free probability — the noncommutative analog of classical independence — to compute traces of matrix products involving analytic functions of $W$ and $X$ in the limit of large dimensions (Mingo and Speicher, 2017). To apply free probability for more general $X$, we require the following assumption:
Assumption 11 (Asymptotic freedom of initial parameters, and input-label pairs).
We assume $W$ and $X$ are asymptotically freely independent. Further, we assume the labels $Y(X)$ are asymptotically free of $W$ (but not $X$).
For example, the labels in the student teacher setups of Adlam et al. (2019), Adlam and Pennington (2020) satisfy this condition.
From Proposition 3, we know that for MSE loss the correlation of NFM/AGOP time derivatives under gradient flow at initialization can be written as:
$\begin{split}&\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)=\,\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\cdot\text{tr}\left((X^{\top}Y\mathcal{K}YX)^{2}\right)^{-1/2}\cdot\text{tr}\left((X^{\top}Y\mathcal{K}^{2}YX)^{2}\right)^{-1/2}.\end{split}$ (7)
In our high-dimensional limit, we expect that, for example, the first term can be given as
$\lim_{n,d,k\to\infty}\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)=\lim_{n,d,k\to\infty}\mathbb{E}_{\theta}\left[\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\right].$ (8)
We can decompose the average as
@@ 4.3 Early time C-NFC dynamics in co-scaling regime
$\begin{split}&\mathbb{E}_{\theta}\left[\text{tr}\left(X^{\top}Y\mathcal{K}YXX^{\top}Y\mathcal{K}^{2}YX\right)\right]=\text{tr}\left(X^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}\right]YXX^{\top}Y\mathbb{E}_{\theta}\left[\mathcal{K}^{2}\right]YX\right)+\text{tr}\left(\text{Cov}\left(X^{\top}Y\mathcal{K}YX,X^{\top}Y\mathcal{K}^{2}YX\right)\right)\end{split}$ (9)
with similar decompositions for the denominator term.
Therefore if the statistics of $\mathcal{K}$ can be understood as a function of $X$, we can compute the correlation in this linear triple-scaling limit. We focus for now on a one-hidden layer quadratic network (similar to the previous section) to avoid the branching of terms that arises in more complicated networks. We provide some additional analysis of the first term in Equation 9 for more complicated architectures in Appendix E.
@@ 4.4 Exact predictions with one hidden layer and quadratic activations
Concretely, we study a neural network $f$ which can be written as $f(x)=a^{\top}(Wx)^{2}$, where $a$ is the readout layer, and the square is element-wise. Let $M^{(4)}_{X|Y}=(X^{\top}YX)^{2}$, and $M^{(2)}_{X}=X^{\top}X$, and $F_{a}=W^{\top}{\rm diag}\left(a^{2}\right)W$. Then, the numerator of Equation 9 can be written as:
@@
We reiterate our assumptions and compute this trace (as well as those for the denominator terms) in Appendix F using standard results from random matrix theory. These calculations show us that the correlation is determined by traces of powers of $M^{(4)}_{X|Y}$ and $M^{(2)}_{X}$ by the calculations in Section F, which are properties of the input-label pairs, and $F_{a}$, which is specific to the architecture and initialization statistics.
@@
To numerically explore the validity of the random matrix theory calculations, we developed a method to generate datasets with different values of $\rho\left(\dot{W}^{\top}\dot{W},\dot{W}^{\top}K\dot{W}\right)$. We construct a random dataset called the alignment reversing dataset, parameterized by a balance parameter $\gamma\in(0,1]$ to adversarially disrupt the NFA near initialization in the regime that width $k$, input dimension $d$, and dataset size $n$ are all equal ($n=k=d=1024$). By Proposition 16, for the aforementioned neural architecture, the expected second derivative of the centered NFM satisfies, $\mathbb{E}\left[\dot{W}^{\top}\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}\right]YX=(X^{\top}YX)^{2}$, while the expected second derivative of the centered AGOP, $\mathbb{E}\left[\dot{W}^{\top}K\dot{W}\right]=X^{\top}Y\mathbb{E}\left[\mathcal{K}^{2}\right]YX$, has an additional component $X^{\top}YX\cdot X^{\top}X\cdot X^{\top}YX$. Our construction exploits this difference in that $X^{\top}X$ becomes adversarially unaligned to $X^{\top}YX$ as the balance parameter decreases.
@@
In our experiment, we sample multiple random datasets with this construction and compute the predicted and observed correlation of the second derivatives of the centered NFC at initialization. For specific details on the construction, see Appendix G.
@@ 5.1 Feature learning and initialization
One factor that modulates the level of feature learning is the scale of the initialization in each layer. In our experiments training networks with unmodified gradient descent, we observe that the centered NFC will increasingly dominate the uncentered NFC with decreasing initialization (third panel, first row, Figure 4). Further, as the centered NFC increases in contribution to the uncentered quantity, the strength of the UC-NFC, and to a lesser extent, the C-NFC increases (Figure 4). The decreases in correspondence between these quantities is also associated with a decrease in the feature quality of the NFM (Appendix I), for the chain monomial task.
For fully-connected networks with homogeneous activation functions, such as ReLU, and no normalization layers, decreasing initialization scale is equivalent to decreasing the scale of the outputs, since we can write $f(Wx)=a^{-p}f(aWx)$ for any scalar $a$ for any homogenous activation $f$. This in turn is equivalent to increasing the scale of the labels. Therefore, decreasing initialization forces the weights to change more in order to fit the labels, leading to more change in $F$ from its initialization. Conceptually, this aligns with the substantial line of empirical and theoretical evidence that increasing initialization scale or output scaling transitions training between the lazy and feature learning regimes (Chizat et al., 2019, Woodworth et al., 2020, Agarwala et al., 2020, Lyu et al., 2023).
@@ 5.2 Speed Limited Optimization
$W^{(\ell)}_{t+1}\leftarrow W^{(\ell)}_{t}-\eta\cdot C_{\ell}\cdot\frac{\nabla_{W^{(\ell)}_{t}}\mathcal{L}}{\|\nabla_{W^{(\ell)}_{t}}\mathcal{L}\|}$
where the hyperparameter $C_{\ell}\geq 0$ controls the amount of learning in layer $\ell$. We expect this rule to increase the strength of the UC-NFC in layers where $C_{\ell}$ is large relative to $C_{m}$ for $m\neq\ell$, as $W^{(\ell)}$ will be forced to change significantly from initialization. By forcing the weights in a particular layer $\ell$ to have fixed learning speeds, these weights will have fixed updates sizes at every epoch, regardless of the loss. We downscale the speeds in other layers $\ell\neq m$ to prevent training instability. As a result, $\left\|W_{0}^{(\ell)}\right\|\left\|W_{t}^{(\ell)}\right\|^{-1}\rightarrow 0$ for large $t$, causing the centered and uncentered NFC to coincide for this layer.
We demonstrate the effects of SLO on the chain-monomial task (Equation (2)). We found that fixing the learning speed to be high in the first layer and low in the remaining layers causes the ratio of the unnormalized C-NFC to the UC-NFC to become close to $1$ across initialization scales (Figure 4). The same result holds for the SVHN dataset (see Appendix M). We note that this intervention can be applied to target underperforming layers and improve generalization in deeper networks (see Appendix K for details).
We observe that both the C-NFC and the UC-NFC become close to $1$ after training with SLO, independent of initialization scale (Appendix I). Further, the quality of the features learned, measured by the similarity of $F$ and $\bar{G}$ to the true EGOP, significantly improve and are more similar to each other with SLO, even with large initialization. In contrast, in standard training the UC-NFA fails to develop with large initialization, as $F$ resembles identity (no feature learning), while $\bar{G}$ only slightly captures the relevant features.
@@ 6 Discussion
Analyzing more general settings
In principle our analysis can be extended to a larger sample size and activation functions with uncentered derivatives. Both of these settings require analyzing more terms in the Taylor expansion of the PTK matrix. This is more complicated than studying the loss, as our most complex calculations require understanding degree $8$ polynomials of the inputs even in the simplest case of a one-hidden layer network (as opposed to degree $4$ to understand the loss). We may also want to understand the case that $n=d^{\ell}$ for integer $\ell$ (i.e. without the logarithmic factor we consider). The appropriate Taylor expansions in these cases are discussed in Misiakiewicz (2022), and will likely require additional structure on the coefficients of the target function $f^{*}$ in the basis of spherical harmonics.
@@
Our theoretical analysis in this work shows that the PTK feature covariance at initialization has a relatively simple structure in terms of the weights and the data. To predict the NFC later in training, we will likely need to account for the change in this matrix. One should be able to predict this development at short times by taking advantage of the fact that eigenvectors of the Hessian change slowly during training (Bao et al., 2023). An alternative approach would be to use a quadratic model for neural network dynamics (Agarwala et al., 2022, Zhu et al., 2022) which can capture dynamics of the empirical NTK (and therefore the PTK). One promising avenue to study the time dynamics of the PTK could be to adapt the results of Wang et al. (2024), which study how the effects of feature learning can propagate to the conjugate kernel for neural networks. It may also be useful to unify our notion of alignment with the observations of Arous et al. (2023), where they find that the weights align with the averaged outer products of the loss gradients.
@@
There are prior works that implement differential learning rates across layers (Howard and Ruder, 2018, Singh et al., 2015), though these differ from differential learning speeds as in SLO, in which the L2-norm of the gradients are fixed, and not the scaling of the gradients. Our intervention is more similar to the LARS (You et al., 2017) and LAMB (You et al., 2019) optimizers that fix learning speeds to decrease training time in ResNet and BERT architectures. However, these optimizers fix learning speeds to the norm of the weight matrices, while our intervention sets learning speeds to free hyperparameters in order to increase feature learning, as measured by the NFC. We also point out that as an additional explanation for the success of SLO in our experiments, the PTK feature covariance may change slowly (or just in a single direction) with this optimizer, because the learning rate is small for later layers, allowing the first layer weights to align with the PTK.
Other works consider interventions that vary the strength of feature learning across layers. The authors of Yang and Hu (2020) down-scale the final layer weights so that the weights in earlier layers must move significantly in $\ell_{\infty}$ norm to fit the data. The SLO intervention is more extreme than this approach in that SLO overrides the natural GD dynamics to exactly set the weight movement rates. In Chizat and Netrapalli (2024) the authors consider adjusting learning rates in each layer so as to force the activation vectors to substantially change across all layers. Instead, SLO sets learning speeds of the weights in each layer so that some weights move much faster than others, depending on the layer where we want to maximize feature learning.
@@
We demonstrate in this work that the dataset (by our construction in Section 4) and optimizer (with SLO in Section 5) play a significant role in the strength of feature learning. Important future work would be to understand other settings where significant empirical differences exist between optimization choices. In particular, analyzing the NFC may clarify the role of gradient batch size and adaptive gradient methods in generalization (Zhu et al., 2023).